# Generating, Reconstructing, and Representing Discrete and Continuous Data: Generalized Diffusion with Learnable Encoding-Decoding

### v1

@@ 4.1 Text
We use BERT-small (Devlin et al., 2019; Bhargava et al., 2021) as the encoder and GPT2-xl (Radford et al., 2019) as the decoder. A warmup training is done to align the encoder and decoder using the bookcorpus dataset (Zhu et al., 2015) without diffusion, followed by the proposed unified DiLED training on the Yelp review dataset (Shen et al., 2017; Li et al., 2018). We compare DiLED with LatentOps (Liu et al., 2023a) and Optimus-DAAE (Shen et al., 2020; Li et al., 2020), using the same architecture and training procedure. We also compare with latest text diffusion models, GENIE (Lin et al., 2023) and AR-Diffusion (Wu et al., 2023). In addition, we compare with fine-tuned GPT2-xl and 20-shot GPT4 (Achiam et al., 2023). For comprehensive evaluation, we study tasks of generation, reconstruction, interpolation, and editing. More details are in §C.1.1. We use BLEU score to measure content preservation, MAUVE (Pillutla et al., 2021) to assess the distribution match between the generated samples and the original data, and perplexity to evaluate text fluency. For text editing with latent arithmetic, transfer accuracy is measured as the geometric mean (Liu et al., 2023a) of the accuracy (by BERT classifiers) and the BLEU score (relative to the input text).
@@
LatentOps 87.6 68.1 0.240 0.50 57.3 32.53 0.697
Optimus-DAAE 86.1 94.1 0.006 0.17 51.0 33.72 0.770
@@
GPT2 - 15.0 0.015 0.32 - - -
GPT4 100 25.7 0.007 0.65 49.3 39.7 0.010
DiLED 92.1 16.4 0.977 0.79 57.1 30.84 0.763
Table 1: Evaluation of Text Reconstruction, Generation (§4.1.1), Latent Vector Arithmetic (§4.1.3), and Interpolation (§4.1.2)
@@ 4.1.1 Text Generation and Reconstruction
Figure 1b and Table 1 show the results. Autoencoding-based models (LatentOps and Optimus-DAAE) exhibit proficiency in reconstruction tasks, primarily because their objectives rely heavily on reconstruction quality. However, these models struggle to generate fluent text that aligns with the real data distribution. This limitation stems from the regularization in latent space designed to adhere to a specific prior distribution, such as a standard Gaussian. The regularization, while theoretically sound, often results in a significant disparity between the model’s latent space and the actual latent distribution observed in real-world scenarios (Yang et al., 2017; Li et al., 2020). The text diffusion models GENIE and AR-Diffusion apply a diffusion process over the token space. Despite the reconstruction-based training objective, these models fall short of both reconstruction and generation. The results highlight the difficulty of standard diffusion modeling for discrete text data. The fine-tuned GPT2 autoregressive model excels in generation fluency but performs poorly in terms of generation diversity and domain match. It also fails for reconstruction. Despite the general capability of GPT4, it fails to capture the domain-specific text characteristics via in-context demonstrations. In comparison, DiLED achieves strong performance across all metrics for both reconstruction and generation. The learned autoencoding ensures effective reconstruction, while the diffusion process introduces dynamic regularization to the latent space, ensuring superior quality in text generation.
Figure 3 provides the training curves, showing that DiLED’s training is more stable and converges faster than other models.
Figure 3: Training curves. DiLED converges fast in a stable way. LatentOps, which depends on text VAE, relies on periodic scheduling of the weight of the prior regularization. The scheduling is common for training text VAE models (Li et al., 2020), and leads to the ups-and-downs in the loss curve, indicating a difficult trade-off between generation and reconstruction capabilities.
@@ 4.1.2 Sentence Interpolation
We set the number of interpolation steps to 10 and evaluated the quality at each step to gauge performance. Particularly, the interpolation results at the midpoint are shown in Table 1. DiLED demonstrates a consistent ability to produce fluent sentences that closely align with the original data distribution. In contrast, baseline models such as GENIE and AR-Diffusion struggled, primarily due to their lack of a semantically coherent latent space for whole sentences. Additionally, we assessed the outcomes using MAUVE and BLEU metrics, with detailed interpolation examples provided in §C.1.2.
@@ 4.2 Image
We adopt UNet (Long et al., 2015) as the encoder and the diffusion-based model in DiffAE (Preechakul et al., 2022)) as the decoder. Following DiffAE, we train our model and then evaluate the reconstruction and generation abilities on the popular FFHQ (Karras et al., 2019), CelebA (Karras et al., 2018), LSUN-Bedroom, and LSUN-Horse (Yu et al., 2015). We use the CelebA-HQ dataset to perform image manipulation tasks. In addition to DiffAE, we compare our model with a number of latest models, including the Latent Diffusion Model (LDM, Rombach et al., 2021)), DDIM (Song et al., 2021a), StyleGAN-XL (Sauer et al., 2022), NVAE (Vahdat & Kautz, 2020b), and Consistency Models (Song et al., 2023). We compare DiLED with these models where they have the existing corresponding checkpoints. Following the common practice, we use FID and reconstruction-FID (rFID) to evaluate the generation quality and reconstruction quality, respectively. To systematically evaluate DiLED, besides the individual evaluations of reconstruction and generation, we mainly perform image interpolation and manipulation4.2.2) tasks, which can reflect the integrated ability. We perform interpolation within the latent space and then reconstruct images from the resulting interpolated representations. Given two latent vectors, ${\mathbf{x}}_{1}^{1}$ and ${\mathbf{x}}_{1}^{2}$, we employ linear interpolation using the formula: $\alpha{\mathbf{x}}_{1}^{1}+(1-\alpha){\mathbf{x}}_{1}^{2}$, where $\alpha$ represents the interpolation ratio and we use $\alpha=0.2,0.4$. This interpolated latent vector is subsequently fed into the decoder to produce the interpolated image. In total, 50k images are used for interpolation and the obtained images are measured against 50k images in the training set to obtain the FID. A detailed image experimental setup can be found in §C.2.1.
@@
The overall performances for generation, reconstruction and interpolation are shown in Figure 4. For all experiments requiring diffusion process, we fix the inference steps to be $T=50$. From the figure, We could observe: (1) DiLED consistently achieves superior or at least comparable performance across all three evaluated tasks and is the best in terms of aggregate performance across the datasets examined. (2) NVAE demonstrates robust capabilities in the tasks of reconstruction and interpolation. However, its performance in the generative task is notably less competitive compared with other models. (3) LDM, while demonstrating commendable performance in reconstruction, falls short in maintaining quality during interpolation. The detailed experimental results are shown in Table 6 (in §C.2.2).
@@ 4.2.2 Image Manipulation
DiLED 13.48 18.43
Table 2: Manipulation.
Following DiffAE, we train a linear classifier, $\mathbf{y}=\mathbf{w}^{\top}\mathbf{x}_{1}+b$, on 70% of the training data for each attribute. This classifier predicts the attribute based on the representation ${\mathbf{x}}_{1}$. To construct the manipulated image representation, we use $\mathbf{x}_{1}^{\prime}=\mathbf{x}_{1}+\epsilon\mathbf{w}$, where $\epsilon$ is a scalar determining the manipulation magnitude. The manipulated representation $\mathbf{x}_{1}^{\prime}$ is then used to generate the corresponding image. We show the examples in Figure 5. We compare ours with DiffAE as we follow its setting and it is also one of the state-of-the-art attribute-based image manipulation methods. Our evaluation focuses on two aspects: 1) Image Quality Post-Manipulation: As shown in Table 2, DiLED consistently yields high-quality images after manipulation. 2) Alignment with Target Class: We test the linear classifier on the remaining 30% of the dataset. In terms of weighted AUC, DiLED’s representation achieves 0.915, closely matching DiffAE’s 0.917. Notably, DiLED surpasses in accuracy, registering 0.893 against DiffAE’s 0.795. A detailed AUC comparison can be found in §C.2.
o
@@ 4.3 Protein Sequences
Following the setup of ReLSO (Castro et al., 2022), a protein autoencoder, we adopt a simple transformer as the encoder and convolutional layers as the decoder. In line with the settings of ReLSO, we jointly train a simple regressor to predict the fitness values from the latent embeddings of protein sequences. This fitness value, representing some desired properties of proteins, serves as a performance metric with higher values indicating superiority. Our models are trained and evaluated on the Gifford (Liu et al., 2019) dataset and GFP (Sarkisyan et al., 2016) dataset separately. In addition to ReLSO, we provided quantitative comparisons between our models and several baseline models, namely NOS (Gruver et al., 2023) which utilized transformer-based Discrete Diffusion and Gaussian Diffusion model for protein design, as well as vanilla Variational Auto-Encoders(Kingma & Welling, 2014). Detailed experimental setups are in §C.3.1. We evaluate DiLED’s representation ability in §4.3.1 and its generation ability through both protein optimization (§4.3.2) and protein generation(§C.3.4). The evaluation on DiLED’s reconstruction ability is provided in the appendix at §C.3.3.
Figure 6: DiLED shows robust performance on fundamental tasks about protein sequences. R, G, and O denote Representation, Generation, and Optimization respectively. Results are normalized for better visualization. Detailed results in §C.3
@@ 4.3.1 Protein Representation
After training, each protein sequence in the test set is transformed into a latent representation, upon which the fitness value is predicted using the regressor in the latent space. Evaluation metrics including MSE, L1 Norm, Pearson and Spearman correlation coefficient are computed between the predicted values and the ground truth; these results are presented in Table 8. The regressor trained in the latent space of DiLED demonstrates superior performance over the regressor from the baseline models on all four metrics on the Gifford Dataset. This suggests that DiLED obtained more refined representations for the protein sequences, leading to more accurate predictions by the regressor. This is also evident from the visualization of the latent space. In Figure 7, we display the latent spaces for both ReLSO(left) and DiLED (right), coloring proteins by their respective fitness value intervals. In the latent space of ReLSO in Fig. 7(left), proteins with fitness values less than 0 exhibits an overlap(inside the red box). While some overlap persists across intervals inside the red box in DiLED’s latent space in Figure 7 (right), the delineation between each interval is more clear.
@@ 4.3.2 Protein Optimization
We optimize a protein sequence by optimizing its corresponding representation in the latent space. We adopt the sampling algorithm introduced in LatentOps that solves an ODE involving the regressor (Liu et al., 2023a). This approach requires a target fitness value to guide the optimization. We set this value to 1, 1.5, 2, and 2.5 for Gifford proteins; 3, 4 for GFP proteins, all of which represent reasonably high fitness values within the respective dataset. Visualizations of the optimized sequence with ReLSO and DiLED algorithms are shown in Figure 8. The grey dots represent proteins from the Gifford dataset, while colored dots represents optimized proteins with varying target fitness values. As shown in Figure 8 (left), the pseudo-convex nature of ReLSO leads to convergence of optimized sequences to a singular point, revealing a lack of diversity. In contrast, as depicted in Figure 8 (right), our model not only achieves superior fitness values but also fosters a broader protein variety. For a detailed quantitative analysis, refer to §C.3.2.
Figure 7: Protein Latent Space of ReLSO (left) and DiLED (right). Different colors represent different fitness values of the corresponding protein sequence.
Figure 8: Optimized Protein Sequence of ReLSO (left) and DiLED (right) in the latent space.
5 Related Work
To tackle the limitations of current generative models, past research has ventured into developing hybrid models by combining VAEs  (Razavi et al., 2019; Vahdat & Kautz, 2020a) and GANs  (Karras et al., 2020a; Esser et al., 2021; Xia et al., 2023a) to create VAE-GAN hybrids  (Larsen et al., 2016; Hu et al., 2018; Wu et al., 2020; Xu et al., 2021). Recently, diffusion models  (Sohl-Dickstein et al., 2015; Ho et al., 2020) have been noted for their generative ability, yet they fall short in providing a robust semantic latent space. Efforts like  Wehenkel & Louppe (2021) and  Vahdat et al. (2021) explored the parameterization of the variational prior with diffusion models. Latent Diffusion Model (LDM,  (Rombach et al., 2021)) and Stable Diffusion  (Rombach et al., 2021) merge autoencoder and diffusion model architectures, optimizing diffusion training with the VAE’s pivotal role in image dimension compression. A deep dive can be found in §D.1. Despite these advancements, it remains elusive to achieve all three capabilities of representation, generation, and reconstruction effectively. DiLED aims to tackle these challenges.

### v2

@@ 4 Experiments
In this section, we present the main experimental results on text, image, and protein data. Additional results and related analysis can be found in §C.
@@ 4.1 Text
We use BERT-small (Devlin et al., 2019; Bhargava et al., 2021) as the encoder and GPT2-xl (Radford et al., 2019) as the decoder. A warmup training is done to align the encoder and decoder using the bookcorpus dataset (Zhu et al., 2015) without diffusion, followed by the proposed unified EDDPM training on the Yelp review dataset (Shen et al., 2017; Li et al., 2018). We compare EDDPM with LatentOps (Liu et al., 2023a) and Optimus-DAAE (Shen et al., 2020; Li et al., 2020), using the same architecture and training procedure. LatentOps is a form of latent diffusion model with successful applications on text data. It uses ordinary differential equations (ODEs), instead of discrete diffusion process as in standard latent diffusion models, on top of a separately trained VAE latent space. Optimus-DAAE is an improved variant of large-scale text VAEs (Li et al., 2020) inspired by (Shen et al., 2020). We also compare with latest text diffusion models, GENIE (Lin et al., 2023) and AR-Diffusion (Wu et al., 2023). In addition, we compare with fine-tuned GPT2-xl and 20-shot GPT4 (Achiam et al., 2023). For comprehensive evaluation, we study tasks of generation, reconstruction, interpolation, and editing. More details are in §C.1.1. We use BLEU score to measure content preservation, MAUVE (Pillutla et al., 2021) to assess the distribution match between the generated samples and the original data, and perplexity to evaluate text fluency. For text editing with latent arithmetic, transfer accuracy is measured as the geometric mean (Liu et al., 2023a) of the accuracy (by BERT classifiers) and the BLEU score (relative to the input text).
@@
LatentOps 87.6 68.1 0.240 0.50 57.3 32.5 0.697
Optimus-DAAE 86.1 94.1 0.006 0.17 51.0 33.7 0.770
@@
GPT2 - 15.0 0.015 0.65 - - -
GPT4 100 25.7 0.007 0.87 49.3 39.7 0.010
EDDPM 92.1 16.4 0.977 0.79 57.1 30.8 0.763
Table 1: Evaluation of Text Reconstruction, Generation (§4.1.1), Latent Vector Arithmetic (§4.1.3), and Interpolation (§4.1.2). The best results are highlighted in bold, and the second-best results are underlined.
@@ 4.1.1 Text Generation and Reconstruction
Figure 1(b) and Table 1 show the results. Autoencoding-based models (LatentOps and Optimus-DAAE) exhibit proficiency in reconstruction tasks, primarily because their objectives rely heavily on reconstruction quality. However, these models struggle to generate fluent text that aligns with the real data distribution. This limitation stems from the regularization in latent space designed to adhere to a specific prior distribution, such as a standard Gaussian. The regularization, while mathematically sound, often results in a significant disparity between the model’s latent space and the actual latent distribution observed in real-world scenarios (Yang et al., 2017; Li et al., 2020). The text diffusion models GENIE and AR-Diffusion apply a diffusion process over the token space. Despite the reconstruction-based training objective, these models fall short of both reconstruction and generation. The results highlight the difficulty of standard diffusion modeling for discrete text data. The fine-tuned GPT2 autoregressive model excels in generation fluency but performs poorly in terms of generation diversity and domain match. It also fails for reconstruction. Despite the general capability of GPT4, it fails to capture the domain-specific text characteristics via in-context demonstrations. In comparison, EDDPM achieves strong performance across all metrics for both reconstruction and generation. The learned autoencoding ensures effective reconstruction, while the diffusion process introduces dynamic regularization to the latent space, ensuring superior quality in text generation.
@@ 4.1.2 Sentence Interpolation
We set the number of interpolation steps to 10 and evaluated the quality at each step to gauge performance. Particularly, the interpolation results at the midpoint are shown in Table 1. EDDPM demonstrates a consistent ability to produce fluent sentences that closely align with the original data distribution. In contrast, baseline models such as GENIE and AR-Diffusion struggled, primarily due to their lack of a semantically coherent latent space for whole sentences. Additionally, we assessed the outcomes using MAUVE and BLEU metrics, with detailed interpolation examples provided in §C.1.2.
@@ 4.1.3 Latent Vector Arithmetic
4.1.4 Training Efficiency
To show the training efficiency of EDDPMs, we measured the training cost on text data. The results shown in Table 2 validate that training EDDPMs is as efficient as training VAEs and variants (e.g., Optimus-DAAE, LatentOps) with the same encoder-decoder architecture. Due to the added diffusion process, the training time of EDDPM for each epoch is  1.3x that of LatentOps and Optimus-DAAE. However, thanks to the stable training process inherited from the well-established DDPM formulation, EDDPMs avoid the different training tricks necessary in VAEs and variants, such as beta-annealing (Bowman et al., 2016), free bits (Kingma et al., 2017), and cyclic annealing schedule (Fu et al., 2019). This allows EDDPMs to converge in fewer steps, leading to similar overall training cost. Figure 3 shows the training curves.
Models Time / Epoch Overall Training Time
LatentOps 13.98 mins 6.98 hrs
Optimus-DAAE 14.49 mins 7.02 hrs
EDDPM 19.35 mins 7.41 hrs
Table 2: Training cost.
Figure 3: Training curves. EDDPM converges fast in a stable way. LatentOps, which depends on text VAE, relies on periodic scheduling of the weight of the prior regularization. The scheduling is common for training text VAE models (Li et al., 2020), and leads to the ups-and-downs in the loss curve, indicating a difficult trade-off between generation and reconstruction capabilities.
@@ 4.2 Image
We adopt UNet (Long et al., 2015) as the encoder and the diffusion-based model in DiffAE (Preechakul et al., 2022)) as the decoder. Following DiffAE, we train our model and then evaluate the reconstruction and generation abilities on the popular FFHQ (Karras et al., 2019), CelebA (Karras et al., 2018), LSUN-Bedroom, and LSUN-Horse (Yu et al., 2015) datasets. We use the CelebA-HQ dataset to perform image manipulation tasks. In addition to DiffAE, we compare our model with a number of latest models, including the Latent Diffusion Model (LDM, Rombach et al., 2021), DDIM (Song et al., 2021a), StyleGAN-XL (Sauer et al., 2022), NVAE (Vahdat & Kautz, 2020b), and Consistency Models (Song et al., 2023). We compare EDDPM with these models where existing corresponding model checkpoints exist. Following the common practice, we use FID and reconstruction-FID (rFID) to evaluate the generation quality and reconstruction quality, respectively. To systematically evaluate EDDPM, besides the individual evaluations of reconstruction and generation, we also perform image interpolation and manipulation4.2.2) tasks, which can reflect the integrated ability. We perform interpolation within the latent space and then reconstruct images from the resulting interpolated representations. Given two latent vectors, ${\mathbf{x}}_{1}^{1}$ and ${\mathbf{x}}_{1}^{2}$, we employ linear interpolation using the formula: $\alpha{\mathbf{x}}_{1}^{1}+(1-\alpha){\mathbf{x}}_{1}^{2}$, where $\alpha$ represents the interpolation ratio and we use $\alpha=0.2$ and $0.4$. This interpolated latent vector is subsequently fed into the decoder to produce the interpolated image. In total, 50k images are used for interpolation and the obtained images are measured against 50k images in the training set to obtain FID score. The detailed image experimental setup is in §C.2.1.
@@
The overall performances for generation, reconstruction and interpolation are shown in Figure 4. For all experiments requiring diffusion process, we fix the inference steps to be $T=50$. From the figure, We could observe: (1) EDDPM consistently achieves superior or at least comparable performance across all three evaluated tasks, and is the best in terms of aggregate performance across the datasets examined; (2) NVAE demonstrates robust capabilities in the tasks of reconstruction and interpolation. However, its performance in the generative task is notably less competitive compared with other models; (3) LDM, while demonstrating commendable performance in reconstruction, falls short in maintaining quality during interpolation. The detailed experimental results are shown in Table 7C.2.2).
@@ 4.2.2 Image Manipulation
EDDPM 13.48 18.43
Table 3: Manipulation (FID).
Following DiffAE, we train a linear classifier, $\mathbf{y}=\mathbf{w}^{\top}\mathbf{x}_{1}+b$, on 70% of the training data for each attribute. This classifier predicts the attribute based on the representation ${\mathbf{x}}_{1}$. To construct the manipulated image representation, we use $\mathbf{x}_{1}^{\prime}=\mathbf{x}_{1}+\epsilon\mathbf{w}$, where $\epsilon$ is a scalar determining the manipulation magnitude. The manipulated representation $\mathbf{x}_{1}^{\prime}$ is then used to generate the corresponding image. We show the examples in Figure 5. We compare ours with DiffAE as we follow its setting and it is also one of the state-of-the-art attribute-based image manipulation methods. Our evaluation focuses on two aspects: 1) image quality post-manipulation: as shown in Table 3, EDDPM consistently yields high-quality images after manipulation. 2) alignment with target class: we test the linear classifier on the remaining 30% of the dataset. In terms of weighted AUC, EDDPM’s representation achieves 0.915, closely matching DiffAE’s 0.917. Notably, EDDPM surpasses in accuracy, registering 0.893 against DiffAE’s 0.795. A detailed AUC comparison can be found in §C.2.
@@ 4.3 Protein Sequences
Following the setup of ReLSO (Castro et al., 2022), a protein autoencoder, we adopt a simple transformer as the encoder and convolutional layers as the decoder. In line with the settings of ReLSO, we jointly train a simple regressor to predict the fitness values from the latent embeddings of protein sequences. This fitness value, representing some desired properties of proteins, serves as a performance metric with higher values indicating superiority. Our models are trained and evaluated on the Gifford (Liu et al., 2019) dataset and GFP (Sarkisyan et al., 2016) dataset separately. In addition to ReLSO, we provided quantitative comparisons between our models and several baseline models, namely NOS (Gruver et al., 2023) which utilized transformer-based discrete diffusion and Gaussian diffusion model for protein design, as well as vanilla VAEs (Kingma & Welling, 2014). Detailed experimental setups are in §C.3.1. We evaluate EDDPM’s representation ability in §4.3.1 and its generation ability through both protein optimization (§4.3.2) and protein generation (§C.3.4). The evaluation on EDDPM’s reconstruction ability is provided in the appendix at §C.3.3.
Figure 6: EDDPM shows robust performance on fundamental tasks about protein sequences. R, G, and O denote Representation, Generation, and Optimization respectively. Results are normalized for better visualization. Detailed results in §C.3
@@ 4.3.1 Protein Representation
After training, each protein sequence in the test set is transformed into a latent representation, upon which the fitness value is predicted using the regressor in the latent space. Evaluation metrics, including MSE, L1 Norm, Pearson and Spearman correlation coefficients, are computed between the predicted values and the ground truth. The results are presented in Table 9. The regressor trained in the latent space of EDDPM demonstrates superior performance over the regressors from the baseline models on all four metrics on the Gifford Dataset. This suggests that EDDPM obtained more refined representations for the protein sequences, leading to more accurate predictions by the regressor. This is also evident from the visualization of the latent space. In Figure 7, we display the latent spaces for both ReLSO (left) and EDDPM (right), coloring proteins by their respective fitness value intervals. In the latent space of ReLSO in Figure 7(left), proteins with fitness values less than 00 exhibits an overlap (inside the red box). While some overlap persists across intervals inside the red box in EDDPM’s latent space in Figure 7 (right), the delineation between each interval is much clearer.
@@ 4.3.2 Protein Optimization
We optimize a protein sequence by optimizing its corresponding representation in the latent space. We adopt the sampling algorithm introduced in LatentOps (Liu et al., 2023a) that solves an ODE involving the regressor (Liu et al., 2023a). This approach requires a target fitness value to guide the optimization. We set this value to 1, 1.5, 2, and 2.5 for Gifford proteins; 3, 4 for GFP proteins, all of which represent reasonably high fitness values within the respective dataset. Visualizations of the optimized sequence with ReLSO and EDDPM algorithms are shown in Figure 8. The grey dots represent proteins from the Gifford dataset, while colored dots represents optimized proteins with varying target fitness values. As shown in Figure 8 (left), the pseudo-convex nature of ReLSO leads to convergence of optimized sequences to a singular point, revealing a lack of diversity. In contrast, as depicted in Figure 8 (right), our model not only achieves superior fitness values but also fosters a broader protein variety. For a detailed quantitative analysis, refer to §C.3.2.
Figure 7: Protein Latent Space of ReLSO (left) and EDDPM (right). Different colors represent different fitness values of the corresponding protein sequence.
Figure 8: Optimized Protein Sequence of ReLSO (left) and EDDPM (right) in the latent space.
5 Other Related Works
To tackle the limitations of current generative models, past research has ventured into developing hybrid models by combining VAEs  (Razavi et al., 2019; Vahdat & Kautz, 2020a; Dai & Wipf, 2019) and GANs  (Karras et al., 2020a; Esser et al., 2021; Xia et al., 2023a) to create VAE-GAN hybrids  (Larsen et al., 2016; Hu et al., 2018; Wu et al., 2020; Xu et al., 2021). Recently, diffusion models  (Sohl-Dickstein et al., 2015; Ho et al., 2020) have been noted for their generative ability, yet they fall short in providing a robust semantic latent space. Efforts like  Wehenkel & Louppe (2021) and  Vahdat et al. (2021) explored the parameterization of the variational prior with diffusion models. Latent diffusion models (e.g., stable diffusion)  (Rombach et al., 2021) merge autoencoder and diffusion model architectures, optimizing diffusion training with the VAE’s pivotal role in image dimension compression. A deep dive can be found in §D.1. Despite these advancements, it remains elusive to achieve all three capabilities of representation, generation, and reconstruction effectively. EDDPMs aim to tackle these challenges.